Check my work Exercise 5-11 Missing Data; Basic CVP Concepts [LO5-1, LO5-9] Fill in the missing amounts in each of the eightCA (For all requirements, Loss amounts should be indicated by a minus sign.) Complete this question by entering your answers(For all requirements, Loss amounts should be indicated by a minus sign.) Complete this question by entering your answers in
Answer #1

Solution:

a)

Case 1 Case 2 Case 3 Case 4
Unit sold 8,400 13,900 20,500 4,000
Sales 218,400 219,900 451,000 123,200
variable expenses 117,600 136,500 205,000 79,200
Fixed expenses 92,000 176,000 81,000 82,000
net operating income(loss) 8,800 (92,000) 165,000 (38,000)
Contribution margin per unit 12 6 12 11

working:

case 1 case 2 case 3 case 4
Unit sold 8,400 =(219,900 -136,500)/6 20,500 4,000
Sales 218,400 219,900 =205,000+165,000+81,000 123,200
Variable expenses 117,600 =219,900 - 176,000 +92,600 205,000 =(123,200 - 82,000 +38,000)
Fixed expenses 92,000 176,000 =(20,500*12)-165,000 82,000
Net operating income (loss) =218,400 - 117,600 -92,000 -92,600 165,000 -38,000
Contribution margin per unit =(218,400-117,600)/8,400 6 12 =(123,200 - 79.200)4000

b)

case 1 case 2 case 3 case4
Sales 441,000 192,000 703,000, 309,000
Variable expenses 255,780 132,480 119,510 92,700
Fixed expenses 111,000 69,000 473,000 313,300
net operating income (loss) 74,220 (9,480) 110,490 (9,700)
Contrbution margin ratio(percentage) 42 31 83 70

Working:

Case 1 Case2 case 3 case 4
Sales 441,000 192,000 =(473,000 + 110,490)/83% 309,000
variable expenses 441,000*58% 132,480 =703,000*17% 92,700
fixed expenses =441,000 - 255,780 - 74,220 69,000 473,000 = 309,000-92,700+97,00
Net oprating income 74,220 =192,000-132,480-69,000 110,490 -9,700
Contrbution margin ratio 42 =(192,000 - 132,480)/192,000 83 =(309,000 - 92,700)/3-9,000